An unknown nonvolatile solute weighing 2.64 g is dissolved in 36.5 g of carbon

tetrachloride, CCl4. The addition of the solute causes an elevation of 1.25o

C in the


boiling point. What is the molar mass of the solute?


1
Expert's answer
2021-01-05T06:44:10-0500

∆Tb = Kbm

Kb for CCl4 = 4.88 °C/mol/Kg

∆Tb =1.25 °C

m = n/W where W = 0.0365Kg (mass of solvent)

so ,

1.25 = 4.88 × (n/0.0365)

On solving n = 0.0093 moles


molar mass of solute = 2.64/0.0093 = 283.8 gram